OpenStudy (anonymous):

The length of a diagonal square is 6 mm. Find the area of the square. A)14.5 B)19.08 C)72 D)18

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Answer is D..

OpenStudy (anonymous):

D I THINK

OpenStudy (anonymous):

If the side length of the square is \(s\), then the diagonal length will be \(d=s\sqrt{2}\) so \(s=\dfrac{d}{\sqrt{2}}\). Thus, the area is \(A=s^2=\left(\dfrac{d}{\sqrt{2}}\right)^2=\boxed{\dfrac{d^2}{2}}\).

OpenStudy (anonymous):

thanks guys